40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2643     Entry time: Fri Feb 26 11:48:36 2010
Author: Koji 
Type: Update 
Category: General 
Subject: MC incident beam shift 

Last night I worked on the MC incident beam such that we can hit the center of the MC mirrors.

Steve and I checked the incident beam on MC1. We found the beam is ~5mm south.
This was not too critical but it is better to be realigned. I moved the steering mirror on the OMC
table (in vac). We kept the MC resonated. After the maximization of the resonance, I realigned the
MC1 and MC3 such that the resonance in dominated by TEM00.

Jenne, Kiwamu, and I then closed the light door on to the OMC/IMC.

I will make more detailed entry with photos in order to explain what and how I did.

ELOG V3.1.3-